OpenStudy (anonymous):

"find all real solutions" 5x^4+29x^3-x^2+29x-6 = 0 HELP PLEASE PRECAL IS SO HARD!

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5x^4+29x^3-x^2+29x-6=\left(5x-1\right)\left(x+6\right)\left(x^2+1\right)}\)

OpenStudy (solomonzelman):

After factoring, you can solve for x by solving three separate cases. *{1}* \(\large\color{black}{ \displaystyle \left(5x-1\right)=0\\[0.5em] }\) *{2}* \(\large\color{black}{ \displaystyle \left(x+6\right)=0\\[0.5em]}\) *{3}* \(\large\color{black}{ \displaystyle \left(x^2+1\right)=0\\[0.5em]}\)

OpenStudy (solomonzelman):

If you are *not* considering any *imaginary* solutions, then case *{3}* will not have any solutions.